A father distributes R1500 to his three children, Jane, Karen and Eddie, in the ratio J:K:E respectively.
If Jane gets twice as much as Karen, who in turn gets three times as much as E . How much does each child receive
1
Expert's answer
2019-05-22T09:31:51-0400

J+K+E=1500

J=2K

K=3E.

So, 2*3E+3E+E=1500

10E=1500, E=1500/10=R150

K=3E=150*3=R450

J=2K=450*2=R900.

Jane received R900, Karen received R450, Eddie received R150.